# M2P2: A Multi-Modal Passive Perception Dataset for Off-Road Mobility in Extreme Low-Light Conditions

### v1

@@ Abstract
Long-duration, off-road, autonomous missions require robots to continuously perceive their surroundings regardless of the ambient lighting conditions. Most existing autonomy systems heavily rely on active sensing, e.g., LiDAR, RADAR, and Time-of-Flight sensors, or use (stereo) visible light imaging sensors, e.g., color cameras, to perceive environment geometry and semantics. In scenarios where fully passive perception is required and lighting conditions are degraded to an extent that visible light cameras fail to perceive, most downstream mobility tasks such as obstacle avoidance become impossible. To address such a challenge, this paper presents a Multi-Modal Passive Perception dataset, M2P2, to enable off-road mobility in low-light to no-light conditions. We design a multi-modal sensor suite including thermal, event, and stereo RGB cameras, GPS, two Inertia Measurement Units (IMUs), as well as a high-resolution LiDAR for ground truth, with a novel multi-sensor calibration procedure that can efficiently transform multi-modal perceptual streams into a common coordinate system. Our 10-hour, 32 km dataset also includes mobility data such as robot odometry and actions and covers well-lit, low-light, and no-light conditions, along with paved, on-trail, and off-trail terrain. Our results demonstrate that off-road mobility is possible through only passive perception in extreme low-light conditions using end-to-end learning and classical planning. The project website can be found at https://cs.gmu.edu/~xiao/Research/M2P2/.
@@ I Introduction
  • a multi-sensor calibration procedure for multi-modal perceptual streams, i.e., infrared radiation caused by temperature differences, per-pixel brightness changes (events), RGB pixels, gyroscope, accelerometer, and magnetometer measurements, and high-resolution point clouds;
@@
  • preliminary results demonstrating that off-road mobility is possible through only passive perception in extreme low-light conditions with end-to-end learning and classical planning methods.
@@ II-A Off-Road Perception
Perception in off-road environments requires both exteroceptive and interoceptive sensing to understand the environment and the robot’s interaction with it. The availability of a wide array of sensors makes safe traversal through off-road environments possible. While a single modality may suffice for navigation in structured environments, the inclusion of multiple modalities in challenging environments adds robustness and redundancy, ensuring that navigation can continue even if one or more sensors are unable to work at full capacity because of adverse environmental conditions. By combining complementary data from multiple sensors, robots can also better perceive and interpret complex environmental features for comprehensive understanding in a variety of off-road unstructured scenarios.
Active sensing modalities like LiDAR and RADAR detect and perceive environmental geometry, enabling the creation of 2D, 3D, or 2.5D elevation maps [8, 9, 10, 11, 12] of the environment. Although LiDAR-based systems are highly popular for their robustness and precision, they can suffer in heavy rain, snow, and fog, and may struggle to map terrain at greater distances [13]. Additionally, the use of pulsated beams can expose the presence of the robot. On the other hand, vision-based navigation systems utilize visible light imaging sensors, e.g., RGB or RGB-D cameras, to understand the terrain semantics [14, 15, 16, 17], create elevation maps [13, 15], and map off-road terrain [18, 19]. Although vision-based navigation systems are advantageous due to their passive sensing capabilities and ability to provide rich environmental information, their reliance on visible light causes poor performance in low-light conditions. While also being passive, interoceptive sensors like IMUs and force sensors measure robot internal states during environment interactions, which can be used to generate traversability maps [14, 20] and model terrain response [21, 22] when combined with exteroception.
@@
II-B Passive Perception
Passive perception sensors detect and interpret environmental data without actively emitting signals. This characteristic makes them particularly useful in situations where stealth is crucial or where active sensors like LiDAR or RADAR might be less effective due to environmental conditions such as heavy rain, fog, or snow.
Stereo RGB Cameras. Stereopsis is a type of passive stereo that is widely used for various applications. This class of approaches compares two images from different viewpoints to estimate scene depth, which is analogous to the binocular vision of human eyes. Challenges like scale drift and failure during pure rotations due to the lack of depth information from a single camera can be effectively solved by using stereo or RGB-D cameras, which enable more reliable visual SLAM solutions [26].
Thermal Camera. Thermal cameras are particularly valuable in scenarios where visible light is insufficient. These sensors detect heat signatures, providing a different perspective on the environment that can complement RGB cameras. In terms of robotics applications, Vidas et al. [27] proposed a method for 3D thermal mapping of building interiors by using a RGB-D and thermal camera, while Aditya et al. [28] used a thermal camera for night-time autonomous driving.
Event Camera. Event camera [29] is an imaging sensor that responds to local changes in brightness to generate spike events. The sensor outputs these events as an asynchronous stream of digital pixel addresses. In recent years, event cameras have been widely used in autonomous vehicles and robotics due to their low motion blur and high dynamic range. These characteristics make them particularly well-suited for challenging off-road environments where camera stabilization is difficult and lighting conditions are poor.
Inertia Measurement Unit (IMU). These sensors monitor the robot’s internal states, such as angular velocities (gyroscope) and translational acceleration (accelerometer), as well as the orientation with respect to the earth’s magnetic field (magnetometer). Such data can be used to assess the robot’s stability, detect slippage, and understand how the terrain responds to the robot’s movements.
II-C Related Datasets
A few existing datasets provide a variety of sensor modalities and ground truth data, enabling the development and benchmarking of algorithms in areas such as SLAM, object recognition, and autonomous navigation: MVSEC [30] is the first dataset that synchronizes stereo event cameras and provides accurate ground truth depth from LiDAR and SLAM and ground truth pose using a motion capture system and GPS; UZH-FPV [31] dataset utilized fast, aggressive, and agile drones to capture event camera data for extreme motion scenarios, but does not contain depth information; For night and day place recognition tasks, Maddern and Vidas [32] built a capture platform consisting of GPS, RGB camera, and thermal camera to capture data from before dawn to after dusk; The KAIST Multi-Spectral Day/Night Dataset [33] introduced a sensor system designed for SLAM, comprising stereo RGB cameras, LiDAR, and thermal camera; Aiming at off-road environments such as forests and urban areas, M3ED [34] used high resolution stereo event cameras, grayscale and RGB cameras, IMU, LiDAR, and RTK localization to collect a high-speed dynamic motion dataset; ViViD++ [25] is the first dataset to feature aligned information from multiple types of alternative vision sensors, including RGB, thermal, event, depth, and inertial measurements. Compared to existing datasets, our M2P2 dataset is the first dataset that focuses on off-road mobility in extremely low-light environments with the most perception modalities and highest sensor quality, as well as a precise multi-modal calibration procedure with accurate synchronization.
@@ III Multi-Modal Sensor Suite
Our multi-modal sensor suite comprises a thermal and an event camera, stereo RGB cameras, two IMUs, GPS, and LiDAR for ground truth. All sensors are assembled on a custom-designed, 3D-printed structure, which can be easily mounted on most mobile robot platforms (Fig. 2). The total dimensions of the sensor suite are 0.31$\times$0.26$\times$0.24 m, with a total weight of 2 kg.
@@ III-C Stereo RGB Cameras
We use two FLIR Blackfly S cameras for capturing images in the RGB spectrum. The cameras have a resolution of 1616$\times$1240, which can be captured at a maximum of 175 FPS (fixed at 10 FPS). While our stereo RGB cameras fail to perceive in no-light conditions, they can still perceive in environments featuring only partial degradation or with some ambient lighting.
@@ III-D IMUs
We use a Yahboom 10-DoF IMU featuring a 3-axis accelerometer, 3-axis gyroscope, 3-axis magnetometer, and a barometer. The sample rate of the IMU is 200 Hz. It features built-in data fusion and gyro stabilization. We also include the IMU embedded in the LiDAR (see details below).
@@ IV Sensor Suite Calibration
Traditional calibration methods use distances measured by geometric features, such as a printed black and white checkerboard with squares of known sizes for camera intrinsics and camera-to-camera extrinsics calibration, or a flat surface for LiDAR-to-camera extrinsics calibration. However, for our multi-modal sensor suite, those methods cause a problem as conventional calibration targets are not visible in the infrared range of a thermal camera. Furthermore, static calibration targets are not visible by an event camera, which needs motion to detect the changes in intensity. Therefore, our multi-modal sensor suite requires a common calibration target that can be perceived by all sensors as to calibrate both intrinsic and extrinsic parameters.
@@ IV-A Thermal Checkerboard
The first challenge of calibrating our sensor suite comes from the thermal camera, which requires different thermal signatures to reflect distances of geometric features. To introduce a contrast thermal signature, we create a calibration target using an aluminum sheet of 3 mm thickness and carbon fiber squares of 35 mm. The sheet and the carbon fiber squares are cut into precise shape using a CNC milling machine achieving an accuracy of 0.05 mm. Since the aluminum sheet reflects most of the long wave infrared (IR) radiation (similar to a mirror in the visible spectrum), we anodize the aluminum sheet to eliminate unwanted reflection in the IR spectrum. After heating the calibration target to roughly 45°$\degree$°C, due to a large difference in emissivity of aluminum and carbon fiber, the checkerboard pattern appears in the thermal image (Fig. 3 left). Due to the contrast in color of aluminum and carbon fiber, the same pattern is visible in both RGB cameras (Fig. 3 right).
@@ IV-C Multi-Modal Synchronization
With a common calibration target visible in all four cameras in the sensor suite (Fig. 3, with another RGB camera in the stereo pair), the last calibration challenge is the precise synchronization among multiple asynchronous and un-synchronized data streams to achieve calibration convergence. To address this, we implement a synchronization scheme as illustrated in Fig. 4. We synchronize all four cameras to the LiDAR, which generates a 10 Hz sync pulse aligned to its encoder angle at 360°. This pulse triggers frame acquisition in the RGB and thermal cameras, with its edges marking temporal points in the event camera stream. The pulse width matches the RGB camera’s exposure time, and its falling edge is used for event camera frame reconstruction. This approach aligns the reconstructed frame with the RGB camera’s exposure completion, ensuring precise temporal correlation across all sensors.
@@ V Multi-Modal Passive Perception Dataset
Our M2P2 dataset encompasses over 10 hours of data collected across various challenging terrain conditions (Fig. 5). The data are gathered with the sensor suite mounted on a Clearpath Husky A200 robot. The dataset includes sequences from a diverse range of environments, progressing from fully prepared paved trails to non-paved off-road paths, and ultimately to unprepared off-trail environments within densely forested areas featuring thick vegetation and narrow passages. To capture a comprehensive range of lighting conditions, data collection is conducted at dusk, with luminosity levels varying from 20 lx to complete darkness (0 lx). This approach ensures the dataset’s applicability to both well-lit and no-light scenarios, addressing the challenges of navigation in varying environmental conditions.
The dataset is structured as ROS-bag files, consisting of compressed RGB and thermal images at 10 FPS, asynchronous raw event stream, 3D point cloud data from LiDAR, IMU data, GPS coordinates, robot odometry and status messages, and human-commanded joystick inputs. All camera data are synchronized using the trigger pulse from the LiDAR, ensuring temporal alignment across multi-modal sensor inputs. Due to the dense canopy of the trees the GPS data is only available for 87.97% of the total dataset. To facilitate accurate sensor placement replication, we provide the URDFs (Unified Robotics Description Format) for the sensor suite configuration on the Husky platform, along with the calibrated transformations. Table I shows the main statistics of the M2P2 dataset. Thanks to our multi-modal synchronization, only six extra RGB images are out of sync with LiDAR point clouds. The slightly fewer number of thermal images is because the thermal driver drops the shutter for about 0.5 s during unexpected ROS node (un)subscription.
TABLE I: M2P2 Statistics
@@
VI Preliminary Results
We conduct two experiments using our M2P2 dataset to demonstrate its usefulness in off-road navigation under degraded lighting conditions.
VI-A End-to-End Learning
To demonstrate the effectiveness of the dataset to enable end-to-end learning for autonomous navigation, we train an end-to-end behavior cloning (BC) model that outputs linear and angular velocities [37, 38] based on thermal camera input into a ResNet-18. Considering the difference in absolute temperature, we normalize each pixel value based on the max and min values of the current thermal image to get relative temperature readings. We deploy this BC model on the Husky robot for a 3.6 km autonomous navigation task on a paved hiking trail, as illustrated in Fig. 7. The luminosity during the experiment ranges from 235 lx to 0 lx (indicated by the color of the path), with the robot completing the majority of the navigation in complete darkness (0 lx). The robot successfully completes the navigation, requiring only 11 human interventions when it goes off-course. Most interventions are because the pavement and the gravel on the side show similar temperature in the thermal input and therefore confuse the robot. More sophisticated techniques that leverage other sensor modalities, e.g., event camera, are necessary to enable more robust navigation.
@@ VI-A End-to-End Learning
VI-B Classical Planning
To showcase the possibility of using fully passive perception in classical navigation approaches, our second experiment utilizes the ROS move_base package for obstacle avoidance. We feed thermal camera images into the DepthAnything [39] model, which provides depth estimations in the form of depth images. These depth images are then converted to laser scan data, which guide the move_base stack in identifying and avoiding obstacles. All computations, including the conversion of thermal images to depth images to laser scans, as well as the navigation itself, are performed on-board the robot. Fig. 8 provides a visual comparison of the thermal image input, the corresponding depth estimation from the DepthAnything model, and the resulting laser scan and costmap used for navigation.
In three out of five classical planning trials, the move_base stack sometimes treats open spaces as obstacles and takes longer detours than necessary, due to a few “ghost obstacles” registered on the costmap from the DepthAnything model output. Despite high quality depth estimation on well-lit RGB images, the DepthAnything model suffers from performance degradation and introduces more obstacle noise when taking night-time thermal images as input. Such a limitation suggests another potential use case of our collected dataset, i.e., fine-tuning existing depth estimation, semantic segmentation, and scene understanding models with the passive perception data and ground truth data in M2P2.
Figure 8: Classical Navigation with Thermal Input. Top Left: Thermal Input; Top Right: DepthAnything Output; Bottom Left: Experiment Lighting Condition; Bottom Right: move_base with Thermal-Produced Scan (white dots).

### v2

@@ Abstract
Long-duration, off-road, autonomous missions require robots to continuously perceive their surroundings regardless of the ambient lighting conditions. Most existing autonomy systems heavily rely on active sensing, e.g., LiDAR, RADAR, and Time-of-Flight sensors, or use (stereo) visible light imaging sensors, e.g., color cameras, to perceive environment geometry and semantics. In scenarios where fully passive perception is required and lighting conditions are degraded to an extent that visible light cameras fail to perceive, most downstream mobility tasks such as obstacle avoidance become impossible. To address such a challenge, this paper presents a Multi-Modal Passive Perception dataset, M2P2, to enable off-road mobility in low-light to no-light conditions. We design a multi-modal sensor suite including thermal, event, and stereo RGB cameras, GPS, two Inertia Measurement Units (IMUs), as well as a high-resolution LiDAR for ground truth, with a multi-sensor calibration procedure that can efficiently transform multi-modal perceptual streams into a common coordinate system. Our 10-hour, 32 km dataset also includes mobility data such as robot odometry and actions and covers well-lit, low-light, and no-light conditions, along with paved, on-trail, and off-trail terrain. Our results demonstrate that off-road mobility and scene understanding under degraded visual environments is possible through only passive perception in extreme low-light conditions. The project website can be found at https://cs.gmu.edu/~xiao/Research/M2P2/.
@@ I Introduction
  • a precise multi-sensor calibration procedure for multi-modal perceptual streams;
@@
  • experimental results demonstrating off-road mobility, depth reconstruction, and vehicle odometry through only passive perception in extreme low-light conditions.
@@ II-A Off-Road Perception
Perception in off-road environments requires both exteroceptive and interoceptive sensing to understand the environment and the robot’s interaction with it. The availability of a wide array of sensors makes safe traversal through off-road environments possible. While a single modality may suffice for navigation in structured environments, the inclusion of multiple modalities in challenging environments adds robustness and redundancy,
ensuring that navigation can continue even if one or more sensors are unable to work at full capacity because of adverse environmental conditions. By combining complementary data from multiple sensors, robots can also better perceive and interpret complex environmental features for comprehensive understanding in a variety of off-road unstructured scenarios.
TABLE I: Comparison with alternative vision datasets.
Sensor Modality
Dataset RGB Depth Thermal Event LiDAR IMU GPS Hardware Environments Lighting
ViViD++ [8] Vehicle Indoor/Urban Day/Night
DiTer++ [9] Legged Diverse Terrain Day/Night
TartanDrive 2.0 [10] Wheeled Off-road Day
M2P2 Wheeled Off-road Day/Night
Active sensing modalities like LiDAR and RADAR detect and perceive environmental geometry, enabling the creation of 2D, 3D, or 2.5D elevation maps [11, 12, 13, 14, 15] of the environment. Although LiDAR-based systems are highly popular for their robustness and precision, they can suffer in heavy rain, snow, and fog, and may struggle to map terrain at greater distances [16]. Additionally, the use of pulsated beams can expose the presence of the robot. On the other hand, vision-based navigation systems utilize visible light imaging sensors, e.g., RGB or RGB-D cameras, to understand the terrain semantics [17, 18, 19, 20], create elevation maps [16, 18], and map off-road terrain [21, 22]. Although vision-based navigation systems are advantageous due to their passive sensing capabilities and ability to provide rich environmental information, their reliance on visible light causes poor performance in low-light conditions.
While also being passive, interoceptive sensors like IMUs and force sensors measure robot internal states during environment interactions, which can be used to generate traversability maps [17, 23] and model terrain response [24, 25] when combined with exteroception.
@@
II-B Related Datasets
A few existing datasets provide a variety of sensor modalities and ground truth data, enabling the development and benchmarking of algorithms in areas such as SLAM, object recognition, and autonomous navigation (Table I): MVSEC [28] is the first dataset that synchronizes stereo event cameras and provides accurate ground truth depth from LiDAR and SLAM and ground truth pose using a motion capture system and GPS; UZH-FPV [29] dataset utilized fast, aggressive, and agile drones to capture event camera data for extreme motion scenarios, but does not contain depth information; For night and day place recognition tasks, Maddern and Vidas [30] built a capture platform consisting of GPS, RGB camera, and thermal camera to capture data from before dawn to after dusk; The KAIST Multi-Spectral Day/Night Dataset [31] introduced a sensor system designed for SLAM, comprising stereo RGB cameras, LiDAR, and thermal camera; Aiming at off-road environments such as forests and urban areas, M3ED [32] used high resolution stereo event cameras, grayscale and RGB cameras, IMU, LiDAR, and RTK localization to collect a high-speed dynamic motion dataset; ViViD++ [8] is the first dataset to feature aligned information from multiple types of alternative vision sensors, including RGB, thermal, event, depth, and inertial measurements. Compared to existing datasets, our M2P2 dataset is the first dataset that focuses on off-road mobility in extremely low-light environments with the most perception modalities and highest sensor quality, as well as a precise multi-modal calibration procedure with accurate synchronization (see Table I for comparison).
@@ III Multi-Modal Sensor Suite
Our multi-modal sensor suite comprises a thermal and an event camera, stereo RGB cameras, two IMUs, GPS, and LiDAR for ground truth. All sensors are assembled in a custom-designed 3D-printed structure, which can be easily mounted on most mobile robot platforms (Fig. 2). The total dimensions of the sensor suite are 0.31$\times$0.26$\times$0.24 m, with a total weight of 2 kg.
@@ III-C Stereo RGB Cameras
We use two FLIR Blackfly S cameras for capturing images in the RGB spectrum. The cameras have a resolution of 1616$\times$1240, which can be captured at a maximum of 175 FPS (fixed at 10 FPS).
While our stereo RGB cameras fail to perceive in no-light conditions, they can still perceive in environments featuring only partial degradation or with some ambient lighting.
@@ III-D IMUs
We use a Yahboom 10-DoF IMU featuring a 3-axis accelerometer, 3-axis gyroscope, 3-axis magnetometer, and a barometer. The sample rate of the IMU is 200 Hz. It features built-in data fusion and gyro stabilization.
We also include the IMU embedded in the LiDAR (see details below).
@@ IV Sensor Suite Calibration
Traditional calibration methods use distances measured by geometric features, such as a printed black and white checkerboard with squares of known sizes for camera intrinsics and camera-to-camera extrinsics calibration, or a flat surface for LiDAR-to-camera extrinsics calibration. However, for our multi-modal sensor suite, those methods pose a limitation as conventional calibration targets are not visible in the infrared range of a thermal camera. Furthermore, static calibration targets are not visible by an event camera, which needs motion to detect the changes in intensity.
Therefore, our multi-modal sensor suite requires a common calibration target that can be perceived by all sensors as to calibrate both intrinsic and extrinsic parameters.
@@ IV-A Thermal Checkerboard
The first challenge of calibrating our sensor suite comes from the thermal camera, which requires different thermal signatures to reflect distances of geometric features. To introduce a contrast thermal signature, we create a calibration target using an aluminum sheet of 3 mm thickness and carbon fiber squares of 35 mm. The sheet and the carbon fiber squares are precision milled with CNC achieving an accuracy of 0.05 mm. Since the aluminum sheet is highly reflective in the long wave infrared (IR) spectrum (similar to a mirror in the visible spectrum), we anodize the aluminum sheet to eliminate unwanted reflection in the IR spectrum. After heating the calibration target to roughly 45°$\degree$°C, due to a large difference in emissivity of aluminum and carbon fiber, the checkerboard pattern appears in the thermal image (Fig. 3 left). Due to the contrast in color of aluminum and carbon fiber, the same pattern is visible in both RGB cameras (Fig. 3 right).
@@ IV-C Multi-Modal Synchronization
With a common calibration target visible in all four cameras in the sensor suite, with another RGB camera in the stereo pair, the last calibration challenge is the precise synchronization among multiple asynchronous and un-synchronized data streams to achieve calibration convergence. To address this, we implement a synchronization scheme as illustrated in Fig. 4.
We synchronize all four cameras to the LiDAR, which generates a 10 Hz sync pulse aligned to its encoder angle at 360°. This pulse triggers frame acquisition in the RGB and thermal cameras, with its edges marking temporal points in the event camera stream. The pulse width matches the RGB camera’s exposure time, and its falling edge is used for event camera frame reconstruction. This approach aligns the reconstructed frame with the RGB camera’s exposure completion, ensuring precise temporal correlation across all sensors.
@@ IV-D All-in-One Calibration Procedure
Fig. 6 shows the LiDAR point cloud overlaid on the corresponding RGB image, along with the reconstructed event frame and thermal image, demonstrating the spatial and temporal alignment of the multi-modal data.
@@
Figure 6: Multi-modal data from the M2P2 dataset, showcasing spatial and temporal alignment in a low-light, off-road forest environment. LiDAR point cloud overlaid on RGB image (left), reconstructed event frame at the trigger’s falling edge (middle), and thermal image (right).
@@ V Multi-Modal Passive Perception Dataset
M2P2 dataset encompasses over 10 hours of data collected across various challenging terrain conditions (Fig. 7). The data are gathered with the sensor suite mounted on a Clearpath Husky A200 robot. The dataset includes sequences from a diverse range of environments, progressing from fully prepared paved trails to non-paved off-road paths, and ultimately to unprepared off-trail environments within densely forested areas featuring thick vegetation and narrow passages. To capture a comprehensive range of lighting conditions, data collection is conducted at dusk, with luminosity levels varying from 20 lx to complete darkness (0 lx). This approach ensures the dataset’s applicability to both well-lit and no-light scenarios, addressing the challenges of navigation in varying environmental conditions.
The dataset is structured as ROS-bag files, consisting of compressed RGB and thermal images at 10 FPS, asynchronous raw event stream, 3D point cloud data from LiDAR, IMU data, GPS coordinates, robot odometry and status messages, and human-commanded joystick inputs. All camera data are synchronized using the trigger pulse from the LiDAR, ensuring temporal alignment across multi-modal sensor inputs. Due to the dense canopy of the trees the GPS data is only available for 87.97% of the total dataset. However, it is possible to fuse LiDAR, IMU, and GPS, when available, with LIO-SAM [35], relying primarily on lidar-inertial odometry. Fig. 8 shows a LIO-SAM-generated map overlaid on a satellite image. The LiDAR point cloud aligns well with visible features (e.g., trail edges and vegetation), demonstrating mapping accuracy. The inset compares the estimated trajectory (blue) to raw GPS (green); the latter deviates significantly under dense tree cover, reflecting degraded signal quality, while the LIO-SAM trajectory remains consistently accurate.
Figure 8: LIO-SAM Mapping Results on Lake Braddock Trail: The LiDAR point cloud (colored points) is overlaid on a satellite image. Inset: Comparison of LIO-SAM estimated trajectory (blue) and raw GPS trajectory (green).
To facilitate accurate sensor placement replication, we provide the URDFs (Unified Robotics Description Format) for the sensor suite configuration on the Husky platform, along with the calibrated transformations. Table II shows the main statistics of the M2P2 dataset. The multi-modal synchronization scheme achieves near-perfect alignment between RGB images and LiDAR point clouds, with only six instances of mis-synchronization. The slightly reduced number of thermal images compared to RGB images is due to the thermal camera’s automatic shutter calibration, which interrupts the image stream for approximately 0.4 seconds to correct for non-uniformities.
TABLE II: M2P2 Statistics
@@
VI Experiment Results
We conduct three experiments using our M2P2 dataset to demonstrate its usefulness in off-road navigation and perception under degraded lighting conditions.
VI-A End-to-End Navigation Learning
To demonstrate the effectiveness of the dataset to enable end-to-end learning for autonomous navigation, we train an end-to-end behavior cloning (BC) model that outputs linear and angular velocities [36, 37] based on thermal camera input into a ResNet-18. Considering the difference in absolute temperature, we normalize each pixel value based on the max and min values of the current thermal image to get the relative temperature readings. We deploy this BC model on the Husky robot for a 3.6 km autonomous navigation task on a paved hiking trail, as illustrated in Fig. 9. The luminosity during the experiment ranges from 235 lx to 0 lx (indicated by the color of the path), with the robot completing most of the navigation in complete darkness (0 lx). The robot successfully completes the navigation, requiring only 11 human interventions when it goes off-course. Most interventions are because the pavement and the gravel on the side show similar temperature in the thermal input and therefore confuse the robot. More sophisticated techniques that leverage other sensor modalities, e.g., event camera, are necessary to enable more robust navigation.
@@ VI-A End-to-End Navigation Learning
VI-B Perception in Degraded Visual Environments
TABLE III: Quantitative Depth Prediction Comparison on Unseen Data.
Model #Params (M) $\downarrow$ Abs Rel $\downarrow$ RMSE $\downarrow$ $\delta_{1}\uparrow$
DepthAnythingV2 335.3 0.66 8.43 0.03
U-Net + M2P2 31 0.13 2.12 0.82
To evaluate the efficacy of M2P2 in enabling scene perception in degraded visual environments, we conduct a comparative analysis of metric depth estimation. Specifically, we train a U-Net [38], 31M parameters, to learn a mapping between thermal infrared imagery and corresponding depth information derived from the LiDAR point clouds. We compare the performance of this U-Net, trained on the M2P2 dataset, against DepthAnythingV2-Large [39], a monocular metric depth estimation model with approximately 335.3 million parameters. Quantitative results, detailed in Table. III, reveal a substantial performance superiority of U-Net despite its significantly lower parameter count. Notably, DepthAnythingV2-Large demonstrates limited generalization to the infrared domain.
Qualitative evaluations, illustrated in Fig. 10, further reinforce these findings. Qualitative inspection confirms that the U-Net trained on M2P2 generates depth maps of considerably higher fidelity compared to those produced by DepthAnythingV2-Large. This observation highlights the pivotal role of domain-specific datasets like M2P2 in enabling the development of robust perception models for degraded visual environments, where traditional RGB-based methods are inherently challenged. Our results suggest that such datasets are indispensable for bridging the gap between standard visual perception and the complexities introduced by atypical sensory inputs.
Figure 10: Qualitative Depth Prediction Comparison on Unseen Data.
VI-C Passive Visual Odometry with Thermal and Event Data
A unique characteristic of M2P2 is the inclusion of calibrated, synchronized thermal and event camera data, enabling exploration of passive perception in extremely low-light conditions. While prior work has investigated visual-inertial odometry using RGB and event cameras [40], the fusion of thermal and event data for odometry remains relatively underexplored. This combination holds significant promise for applications where visible light is scarce or unavailable, such as nighttime off-road navigation or covert operations. The closest existing work is RAMP-VO [40], and M2P2 helps advance this area of research.
To demonstrate the potential of this multi-modal fusion, we adapt the RAMP-VO framework, originally designed for RGB and event data, to process thermal and event data from M2P2. We focus on a challenging 157.5 m segment of the Burke Lake trail to evaluate the robustness of the approach under varying light levels.
Crucially, we simulate reduced lighting conditions by systematically subsampling the event stream. This allows us to assess the performance of the thermal-event odometry system as the available information from the event camera decreases. We experiment with retaining 80%, 50%, and 25% of the original events, representing progressively darker scenarios, in addition to using the full event data (100%).
Table IV presents the translational Absolute Trajectory Error (ATE) for each event subsampling level. As expected, the error generally increases as the event data becomes sparser.
TABLE IV: Translational ATE with Thermal-Event Fusion
Event Percentage Translational ATE (m) $\downarrow$
100% (Full Event Data) 8.79
80% 11.60
50% 12.79
25% 12.49